The coordinates of a kite in the xy plane are x(t)=At and y(t)=3.0m -Bt2 where a=2.4m/s and B=1.2m/s2. Find vectors v(t), a(t) and magnitudes at 2.0s

Possible Answers:

........1: vectors v(t)=A i - 2Bt j, a(t)= -2B j, and magnitude v(2.0s) = -4.8m/s

........2: vectors v(t)=A i - 4Bt j, a(t)= -2B j, and magnitude v(2.0s) = 4.8m/s

........3: None of the above
Answer #1

We have y(t) = 3 - 1.2t^2 an4d x(t) = 2.4t

Plus, we know that v(t) = dx/dt

Hence Vx(t) = 2.4 and Vy(t) = -2.4t

V(t) = 2.4 i - 2.4t j

Further. dV/dt = a(t)

Hence, Ax(t) = 0 and Ay(t) = -2.4

That is magnitude if Velocity at t = 2 seconds is 5.36 m/s

None of the options is correct.